a. A manufacturer’s marginal revenue function is MR=105-x-0.3x^2. Find the increase in
the manufacturer’s total revenue if production is increased from 10 to 20 units.

b. A firm has a marginal revenue given by MR=(3/2x+7)-(1/20), where x is the output. Find
the corresponding demand function.

c. Given that the elasticity of demand for a commodity is given by e_xp=3-2p, where p denotes the price per unit of the commodity. find the demand function x.
1
Expert's answer
2020-07-23T17:45:05-0400

a) "MR = 105 - x - 0.3x^2"

Total Revenue (TR) is found by integrating MR


"TR = \int MR dx"

"= \int (105 - x - 0.3x^2)dx"

"= [105x - \dfrac {x^2} {2} - 0.1x^3 + k]"


Increase in Total revenue is found by TR(20) - TR(10)


"TR(20) = 105(20) - \dfrac {20^2} {2} - 0.1(20)^3 + k"

"= 2100 - 200 - 800 + k"

"= \$(1,100 + k)"

"TR(10) = 105(10) - \dfrac {10^2} {2} - 0.1(10^3) + k"

"= 1050 - 50 - 100 + k"

"= \$(900 + k)"


Thus,

Increase in TR = "\$(1100 + k) - \$(900 + k)"

"= 1100 - 900 + k - k"

"= \$200"




b) "MR = \dfrac {3} {2x + 7} - \dfrac {1} {20}"


"TR = \int MR dx"

"= \int ( \dfrac {3} {2x + 7} - \dfrac {1} {20}) dx"

"= \dfrac {3} {2} ln|2x + 7| - \dfrac {1} {20} x + A"


"AR = \dfrac {TR} {x}"


"= \dfrac {\dfrac {3} {2} ln|2x + 7|} {x} - \dfrac {\dfrac {1} {20} x} {x} + \dfrac {A} {x}"


But, AR = Price (P)

Therefore, the demand function is given by:

"P = = \dfrac {3} {2x} ln|2x + 7| + \dfrac {A} {x} - \dfrac {1} {20}" ,


where A is a constant.

c) "e_{xp} = 3 - 2p"

"\dfrac {dX} {dP} × \dfrac {p} {x} = 3 - 2p"


"\dfrac {dX} {x} = (\dfrac {3} {p} - 2)dP"


"=> \int (\dfrac {1} {x})dX = \int (\dfrac {3} {p} - 2)dP"

"=> ln|x| = 3ln|p| - 2p + lnA"

Where, lnA is a constant

"=> e^{lnx} = e^{lnp^3 - 2p + lnA}"

Therefore, the demand function is given by :


"x = Ap^3e^{-2p}"